OpenStudy (anonymous):

0 = -2x^2 + 8x + 4 Please help find the x intercepts of the graph?

OpenStudy (anonymous):

\[x = 2\pm\sqrt{6}\]

OpenStudy (aravindg):

2+\[\sqrt{6}\]

OpenStudy (aravindg):

2-\[\sqrt{6}\]

OpenStudy (anonymous):

Or what I said.

OpenStudy (aravindg):

:)

OpenStudy (anonymous):

thank youu! :)

OpenStudy (dumbcow):

First factor out a -2 --> -2(x^2 -4x -2) = 0 Then divide by -2 on both sides --> x^2 -4x -2 = 0 Then use quadratic formula to solve for x \[x = \frac{-b \pm \sqrt{b^{2}-4ac}}{2a}\] The solution is whats already been said \[x = 2 \pm \sqrt{6}\]